In the equation x2+mx+k =0, what relation exists between m and k if one of the roots is twice the other?
Answer by stanbon(75887)

In the equation x2+mx+k =0, what relation exists between m and k if one of the roots is twice the other?
--------------
Let one of the roots be "r" and the other root "2r".
Then (x-r)(x-2r)= x^2+mx+k=0
x^2-(3r)x+2r^2 = x^2+mx+k
m=-3r and k=2r^2
============================
Cheers,
Stan H.